# Fast SDE-based Monte Carlo dose calculation for proton therapy validated against Geant4

### v1

@@ 4.3 Inelastic proton-nucleus collision and scatter
where $a(E,u)$ is a known constant given in Section 6.2.2 of osti_1425114. The cdf of this density is invertible, thus the outgoing angle, $\mu$, can be simulated using
@@
with $U\sim\text{Unif}[0,1]$. The setting of compound materials is handled identically to large elastic scattering (Section 4.2.2).
5 Comparison methods
@@
To verify the results from the proposed model, the Monte Carlo simulation toolkit Geant4 (version 11.3) (agostinelli2003geant4) was used as the benchmark. Geant4 was selected for its flexibility, including the ability to enable or disable specific physical processes and customise physics lists. For this study, we have used the prebuilt QGSP_BIC_EMZ physics list, as it provides accurate proton transport modelling via the Binary Cascade model and includes the EMZ option, which offers the most precise electromagnetic physics list in Geant4, optimised for low-energy proton transport.
5.1 Geometry configurations
@@
Both simulations were configured with identical parameters and tested using cubic homogeneous and heterogeneous phantoms. For the homogeneous phantom, a monoenergetic proton beam (energy of 100 or 150 MeV, nozzle radius of 5 mm, energy spread $\sigma_{E}$ of 0.1 keV and radial dispersion $\sigma_{r}$ of 0.05 rad) was directed perpendicularly into water, with no air gap between the source and the phantom. For the heterogeneous phantom, a 100 MeV proton beam with the same beam settings was used to irradiate a cubic phantom with 2 cm of bone, followed by water. The chemical compositions and densities for all materials are the same for Geant4 and the SDE model. For all cases, $1\times 10^{6}$ protons were fired per run, and dose was scored in both cases using a 3D grid of 1 mm3 voxels across the irradiated volume, enabling pointwise comparison of the resulting dose distributions. The SDE model requires angular limits for backscatter and Rutherford events, which were both set at 0.04 rad. Since the SDE model employs a fixed step length, this was set to 0.5 mm for all simulations. For consistency, the maximum step length in Geant4 was constrained to the same value to ensure a fair comparison in terms of time performance.
5.2 Quantitative analysis
For quantitative assessment, we used integral dose-depth curves, central-axis depth-dose distributions, and lateral profiles. For a more rigorous evaluation, a full 3D gamma analysis (Gamma) was performed. The analysis was implemented using the pymedphys Python package, which incorporates methods presented in wendling2007fast. Gamma analysis is a standard method in radiotherapy for comparing dose distributions, as it accounts for both dose differences and spatial discrepancies in a single metric, making it particularly suitable for validating novel dose calculation models.
In this analysis, the dose distribution from the SDE model is treated as the evaluated distribution, while the Geant4 output serves as the reference. For a pair of evaluated and reference points $(\boldsymbol{r_{e}},\boldsymbol{r_{r}})$, the generalised Euclidean distance $\Gamma$ in a combined dose–distance space can be calculated using the dose difference (DD) and distance-to-agreement (DTA) criteria:
$\Gamma(\boldsymbol{r_{e}},\boldsymbol{r_{r}})=\sqrt{\frac{\|\boldsymbol{r_{e}}-\boldsymbol{r_{r}}\|^{2}_{2}}{\text{DTA}^{2}}+\frac{|D_{e}(\boldsymbol{r_{e}})-D_{r}(\boldsymbol{r_{r}})|^{2}}{\text{DD}^{2}}}.$ (15)
Here, $D_{e}(\boldsymbol{r_{e}})$ and $D_{r}(\boldsymbol{r_{r}})$ are the dose values at the evaluated and reference points, respectively, and $\|\boldsymbol{r}\|_{2}$ denotes the usual Euclidean 2-norm of the vector $\boldsymbol{r}$.. The dose difference (DD) and distance-to-agreement (DTA) criteria define the acceptable tolerances for dose and spatial deviation. The DD criterion can be normalised either globally (relative to the maximum dose) or locally (relative to the dose at each reference point). In this study, the local normalisation method was used. A dose threshold of 1% of the maximum dose was applied to exclude low-dose noise, and the gamma analysis was performed with criteria of DD = 1% and DTA = 2 mm.
@@ 5.2 Quantitative analysis
The gamma index $\gamma$ for a given reference point $\boldsymbol{r_{r}}$ is determined by searching over all evaluated points $\boldsymbol{r_{e}}$ within a radius equal to the DTA. To reduce the influence of the discrete voxel grid, the gamma analysis algorithm in pymedphys employs on-the-fly linear interpolation between voxels, allowing $\gamma$ to be evaluated at sub-voxel positions. This process is controlled by the interpolation fraction, which specifies the interpolation step size as a fraction of the DTA criterion. In this study, the interpolation fraction was set to 10, corresponding to a step size of 0.2 mm. This interpolation does not alter the underlying dose distributions, but enables a more accurate determination of the gamma index, which satisfies:
$\gamma(\boldsymbol{r_{r}})=\min_{\boldsymbol{r_{e}}:\|\boldsymbol{r_{e}}-\boldsymbol{r_{r}}\|_{2}\leq\text{DTA}}\{\Gamma(\boldsymbol{r_{e}},\boldsymbol{r_{r}})\}.$ (16)
A point is considered to pass if $\gamma(\boldsymbol{r_{r}})<1$ and fail otherwise. The overall agreement is expressed as the percentage of reference points that satisfy the gamma criterion. Pass rates are strongly influenced by the chosen DD/DTA criteria and the normalisation method. As a benchmark, universal tolerance limits recommended in miften2018tolerance specify a pass rate of $\geq$ 95% using 3%/2 mm criteria with a 10% dose threshold under global normalisation.
6 Results and discussion
6.1 Proton transport in homogeneous medium
This first test involves the irradiation of a cubic water phantom with a size of $\times20$$\times$20 cm3, subdivided into 1 mm3 cubic voxels in both the SDE and Geant4 models. From the 3D dose arrays, a central-axis 2D slice was extracted to visualise the overall dose distributions, which are shown in Figure 3 using a 100 MeV beam. The dose maps, plotted on a logarithmic scale, illustrate the strong agreement between the SDE and Geant4, including in low-dose regions where the SDE accurately captures lateral spread. The main differences are observed in the lowest dose values, corresponding to non-local dose deposits present only in the Geant4 maps. These arise from secondary particles, such as gammas and neutrons, which travel away from the primary beam and interact with atoms at distant locations, producing small, remote dose contributions that the current SDE model does not simulate.
Figure 3: 2D central-axis dose distribution comparison using a monoenergetic 100 MeV proton beam in a homogeneous water phantom.
Based on these findings, a more detailed comparative analysis was performed, including results from a 150 MeV proton beam to assess the SDE output at higher proton energies. The depth-dose curves shown in Figure 4(a) were obtained by integrating the 3D dose distribution over the cross sectional area to obtain a 1D Bragg peak. Deviations between the integrated depth-dose curves are below 3% for a 100 MeV beam and below 6.5% for a 150 MeV when evaluated pointwise before the Bragg peak. Beyond the peak, two factors lead to systematic differences between the models. First, the SDE model applies an energy threshold below which proton transport is terminated, which causes the dose to fall off slightly faster than in Geant4. Second, because the SDE does not model dose contributions from secondary particles such as neutrons and gammas, it does not reproduce the remanent dose at remote locations predicted by Geant4. Together, these effects result in a steeper fall-off and zero dose deposition at greater depths in the SDE results. Nevertheless, the proton range is accurately reproduced, with R90 agreeing within less than 0.1 mm for 100 MeV and within 0.4 mm for 150 MeV. This evidences the accuracy in the electromagnetic processes modelled in our approach. Moreover, the percentage differences in the integrated depth-dose curves are highest at mid-depths and decrease towards the Bragg peak. This behaviour indicates that the SDE model achieves its best agreement with Geant4 in the high-dose region near the Bragg peak, where dose accuracy is most critical.
For a more detailed comparison of the core beam behaviour, differential depth-dose curves were obtained from the central voxels of the 3D dose arrays and are shown in Figure 4(b). These curves are highly sensitive to local variations in scattering, energy loss and voxel sampling, which results in a higher overall percentage difference compared to integrated depth-dose curves, while remaining within 10% for 100 MeV and 20% for 150 MeV. Moreover, these curves are sensitive to the angular limits set in the SDE model for Rutherford and backscatter events. The agreement is highest at shallow depths, after which it starts decreasing. Overall, the trends suggests that the SDE model reproduces the core beam behaviour reliably, while minor discrepancies at depth are expected due to the differences in scattering modelling.
(a) Integral depth-dose comparison.
(b) Central-axis depth-dose distribution comparison.
@@ 6.1 Proton transport in homogeneous medium
Figure 4: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a homogeneous water phantom, including pointwise calculations for percentage dose differences in the lower subplots.
@@
To further evaluate the beam’s lateral spread and its consistency with depth, Figures 5(a) and 5(b) show the lateral profiles at different depths for 100 and 150 MeV, respectively. The SDE model consistently reproduces the overall Gaussian beam shape, showing similar widths to Geant4 across all depths. This indicates that multiple scattering is modelled with high accuracy. Small, systematic overestimation of the absolute dose is confirmed in all profiles, as can be expected because Geant4 models many energy-dissipating processes neglected by the SDE, most notably a larger range of secondary particles. Despite this, the close match in profile widths demonstrates that the spatial dispersion of the beam is well captured, confirming the accuracy of the model in describing lateral transport.
(a) Profiles using 100 MeV proton beam at different depths.
(b) Profiles using a 150 MeV proton beam at different depths.
@@
Figure 5: Lateral profile comparison between SDE (marker points) and Geant4 (solid lines) for two proton energies in a homogeneous water phantom.
To illustrate the potential of our model for clinical purposes, gamma analysis was performed in 3D for thorough comparison for the two proton energies under study. Figure 6 shows central-axis slices from the resulting 3D gamma index matrices for each proton energy under study, illustrating that the gamma indices remain well below unity throughout the high-dose region, confirming strong spatial and dosimetric agreement between the SDE model and Geant4. The largest discrepancies are confined to low-dose regions near the beam periphery, where secondary-particle effects, absent in the SDE model, contribute in Geant4. In addition, Table 1 presents the gamma pass rates for several commonly used criteria. Pass rates exceed 98% for all evaluated criteria, reaching 100% under the conventional 3%/2 mm condition for a 10% dose threshold. These results demonstrate that the SDE model achieves clinically acceptable agreement with Geant4.
(a) Gamma distribution using 100 MeV proton beam.
(b) Gamma distribution using 150 MeV proton beam.
Figure 6: Central-axis gamma index matrices for two proton beam energies in homogeneous water phantom.
Energy (MeV) DD (%) DTA (mm) Threshold (%) Norm. Method Pass rate (%)
100 3 2 10 global 100
3 2 10 local 100
3 2 1 local 98.80
2 1 1 local 98.17
150 3 2 10 global 100
3 2 10 local 100
3 2 1 local 99.60
2 1 1 local 99.26
Table 1: Gamma analysis pass rates in homogeneous water phantom using two monoenergetic proton energies.
6.2 Comparison against different Geant4 physics lists
@@
To contextualise the observed differences between the SDE model and Geant4, Figure 7 compares integral depth-dose curves obtained with different Geant4 physics lists that are commonly used for proton therapy calculations (QGSP_BIC_EMZ, QGSP_BIC_EMY and QGSP_BERT). Comparisons are shown for 100 MeV (Fig. 7(a)) and 150 MeV (Fig. 7(b)) proton beams, with the percentage differences calculated with respect to the reference physics list, QGSP_BIC_EMZ. The SDE curve is also included for reference, illustrating that the deviation between the SDE model and the reference Geant4 configuration lies within the typical range of variability observed among Geant4 physics lists. Notably, while the SDE model exhibits the largest discrepancies in the low-dose tail of the Bragg curve, agreement in the high-dose region near the peak remains within the inter-list variations of Geant4. These results reinforce the physical consistency of the SDE model, while future efforts will focus on reducing these discrepancies through the implementation of secondary particle modelling.
(a) Depth-dose curves using a monoenergetic 100 MeV beam.
(b) Depth-dose curves using a monoenergetic 150 MeV beam.
@@
Figure 7: Integral depth-dose comparison between different Geant4 physical models and the SDE model. The reference physics list (QGSP_BIC_EMZ) is shown in dotted lines.
6.3 Proton transport in heterogeneous medium

### v2

@@ 2.1.5 Parameterisation for inelastic proton-nucleus interactions
where $a(E,u)$ is a known constant given in Section 6.2.3 of Trkov and Brown (2018). The cdf of this density is invertible, thus the outgoing angle, $\mu$, can be simulated using
@@
with $U\sim\text{Unif}[0,1]$. The setting of compound materials is handled identically to large elastic scattering (see Section 2.1.4).
2.2 Benchmark and comparison
2.2.1 Reference Monte Carlo simulations
@@
To verify the results from the proposed model, the Monte Carlo simulation toolkit Geant4 (version 11.3) (Agostinelli et al., 2003) was used as the benchmark. Geant4 was selected for its flexibility, including the ability to enable or disable specific physical processes and customise physics lists. For this study, we have used the prebuilt QGSP_BIC_EMZ physics list, as it provides accurate proton transport modelling via the Binary Cascade model and includes the EMZ option, which offers the most precise electromagnetic physics list in Geant4, optimised for low-energy proton transport.
2.2.2 Test cases
All simulations were configured with identical parameters and evaluated using three phantom geometries designed to probe different dosimetric effects. The first geometry is a homogeneous $\times20$$\times$20 cm3 water phantom, used as a reference case for baseline dose and range in a uniform medium. The second is a slab phantom of the same size, composed of 2 cm water, 1 cm bone, 2 cm compressed lung, and a distal water region. The layers are laterally homogeneous and are intended to assess the impact of longitudinal heterogeneities on proton range and dose deposition. Slab thicknesses were chosen such that all materials are traversed by both 100 MeV and 150 MeV proton beams, while ensuring that the Bragg peak occurs within the final water layer. The final geometry is a water phantom of the same dimensions, containing a 2 cm thick bone insert located in the upper half of the phantom at a depth of 3 cm, designed to investigate the effect of lateral heterogeneities. These configurations are hereafter referred to as the water, slab, and insert phantoms, respectively.
@@ 2.2.2 Test cases
For all cases, a monoenergetic proton beam (energy of 100 or 150 MeV, nozzle radius of 5 mm, energy spread $\sigma_{E}$ of 0.1 keV and radial dispersion $\sigma_{r}$ of 0.05 rad) was directed perpendicularly into water, with no air gap between the source and the phantom. The chemical compositions and densities for all materials are the same for Geant4 and the SDE model. For all cases, $1\times 10^{6}$ protons were fired per run, and dose was scored in both cases using a 3D grid of 1 mm3 voxels across the irradiated volume, enabling pointwise comparison of the resulting dose distributions. The SDE model requires angular limits for backscatter and Rutherford events, which were set at 0.04 and 0.02 rad, respectively. Since the SDE model employs a fixed step length, this was set to 0.5 mm for all simulations. For consistency, the maximum step length in Geant4 was constrained to the same value to ensure a fair comparison in terms of time performance. For both simulators, the energy deposition is assigned at the step midpoint to minimise bias in regions of steep gradients.
2.2.3 Comparison metrics
For quantitative assessment, integral dose-depth curves, central-axis depth-dose distributions, range calculations and lateral profiles were evaluated. In addition, a full 3D gamma analysis (Low, 2010) and 3D voxel-to-voxel dose difference maps were used to quantify spatial agreement between the SDE and Geant4 dose distributions, where the SDE output is treated as the evaluated distribution, while the Geant4 distribution serves as the reference.
For range estimation, the integrated depth–dose curve was linearly interpolated to determine the depth (to 0.1 mm resolution) at which the dose on the distal fall-off decreases to 90% of its maximum (R90). Voxel-wise dose differences between the evaluated and reference distributions were computed within a reference-defined region of interest. The region of interest was defined as voxels receiving at least 1% of the maximum reference dose. Differences were normalised to the 99th percentile of the reference dose, D99, computed within this region, in order to provide a near-maximum dose scale while reducing sensitivity to isolated high-dose outliers.
Gamma analysis was implemented using the pymedphys Python package, which incorporates methods presented in Wendling et al. (2007). Given that gamma analysis is employed here as a numerical similarity metric between two simulated distributions, rather than as a clinical QA criterion, the distance-to-agreement (DTA) parameter is not intended to compensate for setup or alignment uncertainties, and a stricter dose difference (DD) criterion is chosen. In this study, local dose normalisation was applied, with a dose threshold of 1% of the maximum dose to suppress low-dose noise. Given the exact geometric correspondence and a voxel size of 1 mm, strict acceptance criteria of DD = 2% and DTA = 0.5 mm were used throughout. For context, clinical QA guidelines typically report $\geq$ 95% pass rates using 3%/2 mm criteria with a 10% dose threshold under global normalisation (Miften et al., 2018).
@@ 2.2.3 Comparison metrics
To reduce the influence of the discrete voxel grid, the gamma analysis algorithm in pymedphys employs on-the-fly linear interpolation between voxels, allowing $\gamma$ to be evaluated at sub-voxel positions. This process is controlled by the interpolation fraction, which specifies the interpolation step size as a fraction of the DTA criterion. In this study, the interpolation fraction was set to 10, corresponding to a step size of 0.2 mm. This interpolation does not alter the underlying dose distributions, but enables a more accurate determination of the gamma index.
In addition to dosimetric accuracy, computational performance was recorded as a reference metric. Wall-clock runtimes correspond to the elapsed execution time reported by the Unix time utility. All simulations were executed on a MacBook Pro equipped with an Apple M2 Max processor using $1\times 10^{6}$ primary protons in single-thread mode. The reported timings are not intended as a direct performance comparison with clinically optimised GPU-based Monte Carlo engines, but to illustrate the computational cost of the proposed SDE formulation relative to a standard CPU-based Monte Carlo implementation under identical conditions. Due to its structure, the SDE solver exhibits a predictable per-particle execution cost and is well suited to parallel execution, making it a natural candidate for multi-threaded and GPU-based implementations.
3 Results
3.1 Homogeneous water phantom
We present the dose comparison calculations for the baseline case of a water phantom being irradiated with a proton beam of 100 MeV and 150 MeV. This case serves as a foundation for the overall accuracy of the model. From the 3D dose arrays that are obtained in both reference and evaluated models, the integrated depth-dose curves were obtained and shown in Figure 2(a), while differential depth-dose curves were obtained from the central voxels of the 3D dose arrays for a more detailed comparison of the core beam behaviour are shown in Figure 2(b).
@@ 3.1 Homogeneous water phantom
Figure 2: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a homogeneous water phantom, including pointwise calculations for percentage dose differences in the lower subplots. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves.
Deviations between the integrated depth-dose curves are below 3.5% for a 100 MeV beam and below 6.5% for a 150 MeV when evaluated pointwise before the Bragg peak. Beyond the peak, two factors lead to systematic differences between the models. First, the SDE model applies an energy threshold below which proton transport is terminated, which causes the dose to fall off slightly faster than in Geant4. Second, because the SDE does not model dose contributions from secondary particles such as neutrons and gammas, it does not reproduce the remanent dose at remote locations predicted by Geant4. Together, these effects result in a steeper fall-off and zero dose deposition at greater depths in the SDE results. Nevertheless, the proton range is accurately reproduced, with R90 calculations agreeing within less than 0.1 mm for 100 MeV and within 0.6 mm for 150 MeV. This evidences the accuracy in the electromagnetic processes modelled in our approach. Moreover, the percentage differences in the integrated depth-dose curves are highest at mid-depths and decrease towards the Bragg peak. This behaviour indicates that the SDE model achieves its best agreement with Geant4 in the high-dose region near the Bragg peak, where dose accuracy is most critical.
The central-axis depth dose curves are highly sensitive to local variations in scattering, energy loss and voxel sampling, which results in a slightly higher overall percentage difference compared to integrated depth-dose curves, but still remaining within 6% for 100 MeV and 16% for 150 MeV. Moreover, these curves are sensitive to the angular limits set in the SDE model for Rutherford and backscatter events. The agreement is highest at shallow depths, after which it starts decreasing. Overall, the trends suggests that the SDE model reproduces the core beam behaviour reliably, while minor discrepancies at depth are expected due to the differences in scattering modelling.
@@
To further evaluate the beam’s lateral spread and its consistency with depth, Figures 3(a) and 3(b) show lateral profiles at selected depths for 100 and 150 MeV proton beams, respectively. The SDE model consistently reproduces the overall Gaussian beam shape, showing similar widths to Geant4 across all depths. This indicates that the dominant multiple scattering behaviour of the beam is well captured.
@@
Figure 3: Lateral profile comparison between SDE (marker points) and Geant4 (solid lines) for two proton energies in a homogeneous water phantom (a) using a 100 MeV beam, (b) using a 150 MeV beam.
Figure 4: 2D dose distribution comparison between the SDE model and Geant4 in a homogeneous water phantom for monoenergetic proton beams. (a) Central-axis dose distribution and (b) corresponding relative dose difference map for a 100 MeV beam, normalised to the 99th percentile (D99) of the Geant4 reference dose. (c,d) Same quantities for a 150 MeV beam.
Small, systematic differences in the absolute dose are observed across the profiles. Although both the SDE model and Geant4 model scattering processes using partitioning between single-scattering and multiple scattering regimes using angular cutoffs, differences in how these regimes are implemented and coupled to energy deposition lead to modest variations in the proton path-length distribution and the associated dose. These effects exhibit increased sensitivity at lower beam energies, where scattering processes play a larger role. Despite these differences, the close agreement in profile widths demonstrates that the spatial dispersion of the beam is accurately reproduced. A more noticeable deviation is observed at the Bragg peak location for the 100 MeV beam ($x=7.5$ cm), where the peak dose predicted by our approach differs from Geant4. This reflects the sensitivity of the peak region to discretisation and the treatment of angular scattering.
In addition, a central-axis 2D slice was extracted to visualise the overall dose distributions, along with a voxel-wise difference map for the same slice. These results are shown in Figure 4. The dose maps using a 100 and 150 MeV are shown in Figures 4(a) and 4(c), respectively, and illustrate the strong agreement between the SDE and Geant4, including in low-dose regions where the SDE accurately captures lateral spread. The main differences are observed in the lowest dose values, corresponding to non-local dose deposits present only in the Geant4 maps. These arise from secondary particles that interact with atoms at distant locations, producing small, remote dose contributions that the current SDE model does not simulate.
The relative dose difference maps in Figures 4(b) and 4(d) provide a spatially resolved, qualitative assessment of the agreement between the SDE model and Geant4. Since the differences are normalised to a global high-dose reference (D99 of the Geant4 dose), the reported values should not be interpreted as local voxel-wise dose errors. Instead, they indicate the magnitude and spatial distribution of discrepancies relative to the characteristic high-dose scale of the problem. Bearing this in mind, there is good agreement across the high-dose region, with the largest relative differences localised around the distal fall-off, where steep dose gradients amplify small discrepancies between the two models. In the entrance and plateau regions, differences remain low. The pattern and magnitudes of the difference maps is similar for both beam energies, suggesting consistent model behaviour over the energy range.
For completeness and to illustrate the potential of our model for clinical purposes, gamma analysis was performed in 3D for thorough comparison for the two proton energies under study. Pass rates of 97.1% for the 100 MeV case and 95.4% for the 150 MeV case were obtained using stricter criteria than the conventional clinical ones (DD = 2%/0.5 mm, 1% dose threshold). This demonstrates that the SDE model achieves clinically acceptable agreement with Geant4. In terms of computational performance, the SDE model is 2.5 times faster than the single-threaded Geant4 model when using the 100 MeV beam, and 2.8 times faster using the 150 MeV beam.
3.1.1 Comparison against different Geant4 physics lists
@@
To contextualise the observed differences between the SDE model and Geant4, Figure 5 compares integral depth-dose curves obtained with different Geant4 physics lists that are commonly used for proton therapy calculations (QGSP_BIC_EMZ, QGSP_BIC_EMY and QGSP_BERT). Comparisons are shown for 100 MeV (Fig. 5(a)) and 150 MeV (Fig. 5(b)) proton beams, with the percentage differences calculated with respect to the reference physics list, QGSP_BIC_EMZ. The SDE curve is also included for reference, illustrating that the deviation between the SDE model and the reference Geant4 configuration lies within the typical range of variability observed among Geant4 physics lists.
@@
Figure 5: Integral depth-dose comparison between different Geant4 physical models and the SDE model for two beam energies (a) 100 MeV and (b) 150 MeV. The reference physics list (QGSP_BIC_EMZ) is shown in dotted lines.
Notably, while the SDE model exhibits the largest discrepancies in the low-dose tail of the Bragg curve, agreement in the high-dose region near the peak remains within the inter-list variations of Geant4. These results reinforce the physical consistency of the SDE model, while future efforts will focus on reducing these discrepancies through the implementation of secondary particle modelling. Since QGSP_BIC_EMY and QGSP_BERT are computationally less expensive than QGSP_BIC_EMZ, it is most meaningful to compare their runtimes with the SDE model. For this geometry, the SDE implementation is from 2.7 to 3.1 times faster than QGSP_BIC_EMY and from 2.3 to 2.5 times faster than QGSP_BERT.
3.2 Longitudinally heterogeneous slab phantom